OpenStudy (mtalhahassan2):

The area of a triangle can be calculated from Heron's formula. A= square s(s-a)(s-b)(s-c), where a,b,and c are the side lengths and s= a+b+c divide by 2. Calculate the area of a triangle with side lengths 5,7, and 10. Leave your answer in simplest radical form.

OpenStudy (mayankdevnani):

first,find out semi-perimeter of triangle

OpenStudy (mayankdevnani):

\[\large \bf s=\frac{a+b+c}{2}=\frac{5+7+10}{2}=11\]

OpenStudy (mtalhahassan2):

ok

OpenStudy (mayankdevnani):

now,use HERON'S FORMULA :- \[\huge \bf \sqrt{s(s-a)(s-b)(s-c)}\] \[\large \bf \sqrt{11(11-5)(11-7)(11-10)}\]

OpenStudy (mtalhahassan2):

Then what

OpenStudy (mayankdevnani):

find out the value !!

OpenStudy (mayankdevnani):

i mean simplify it and get your answer :)

OpenStudy (mtalhahassan2):

ok

OpenStudy (mayankdevnani):

so what is the answer ???

OpenStudy (mtalhahassan2):

wait a second

OpenStudy (mtalhahassan2):

give me some time

OpenStudy (mayankdevnani):

okk

OpenStudy (mtalhahassan2):

So it be like 2 sqaure 66

OpenStudy (mtalhahassan2):



OpenStudy (mayankdevnani):

correct !!! \[\large \bf 2\sqrt{66}=16.24=\color{red}{Answer}\]

OpenStudy (mayankdevnani):

good job :) @MTALHAHASSAN2

OpenStudy (mtalhahassan2):

Thanks

OpenStudy (mayankdevnani):

your welcome :)